# Negotiative Alignment: Embracing Disagreement to Achieve Fairer Outcomes – Insights from Urban Studies

### v1

@@ Abstract.
Based on these empirical insights, we propose negotiative alignment, an AI framework that treats disagreement as an essential input to be preserved, analyzed, and addressed. Negotiative alignment builds on pluralistic models by dynamically updating stakeholder preferences through multi-agent negotiation mechanisms, ensuring no single perspective is marginalized. We outline how this framework can be integrated into urban analytics—and other decision-making contexts—to retain minority viewpoints, adapt to changing stakeholder concerns, and enhance fairness and accountability. The study demonstrates that preserving and engaging with disagreement, rather than striving for an artificial consensus, can produce more equitable and responsive AI-driven outcomes in urban design.
@@
Negotiative alignment, urban design, AI alignment, multi-stakeholder preferences, fairness, accountability, participatory, inclusivity
ccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Continuous space searchccs: Computing methodologies Theory of computation for AI alignmentccs: Computing methodologies Reinforcement learningccs: Computing methodologies Machine learning approachesccs: Information systems Decision support systemsccs: Applied computing Computational social sciencesccs: Applied computing Collaborative and participatory computingcopyright: noneconference: -; -booktitle: ;
1. Introduction
@@
Urban environments appear to serve everyone, yet people experience cities through different social, cultural, and political lenses. A wheelchair user in Montreal may perceive certain sidewalks as insurmountable, while an LGBTQIA2+ resident might interpret street murals or local businesses as signaling inclusion or exclusion. However, contemporary urban analysis tends to emphasize aggregated metrics: a single “overall rating” for aesthetic quality, accessibility, or safety (Mehta, 2014; Varna and Tiesdell, 2010). Although helpful for city-wide planning, these averages may dilute the specific, and often urgent, needs of marginalized groups (Fainstein, 2010; Costanza-Chock, 2020).
@@
Drawing on a multi-phase study of public space in Montreal, we show that apparent disagreements in urban evaluations are not merely noisy outliers. Rather, they reflect differences in lived experience, structural inequality, and cultural values. Participants in our study included 35 individuals from multiple demographic categories, such as wheelchair users, seniors, LGBTQIA2+ individuals, and immigrants. By combining semi-structured interviews, rating tasks, and forced-ranking experiments, we found that disagreements were pervasive and persistent, particularly concerning dimensions like accessibility, inclusivity, and practicality. As discussed in Section 4, such divergent viewpoints could not be resolved merely by factual clarifications; instead, they pointed to deeper identity-based needs.
@@
Conventional AI or survey-based methods prioritize consensus by aggregating heterogeneous preferences into a single output score. This approach often obscures critical minority viewpoints (Gabriel, 2020; Rawls, 1999), such as the need for comprehensive curb cuts for wheelchair users or the importance of affirming cultural symbols for LGBTQIA2+ communities (Anttiroiko and de Jong, 2020). While recent advances in AI alignment incorporate diverse perspectives through multi-agent modeling, game-theoretic approaches, and fairness-aware optimization (Zhi-Xuan et al., 2024; Sorensen et al., 2024; Jain et al., 2024; Dwork et al., 2012; Ferrara, 2023), they rarely include explicit mechanisms for negotiating conflicting preferences in real time. Addressing this gap requires alignment strategies that iteratively adjust stakeholder weights and trade-offs rather than statically aggregating preferences.
@@
To bridge this gap, we introduce negotiative alignment, a dynamic, multi-agent approach that integrates disagreements into iterative AI decision-making processes (Section 5). Unlike singular consensus methods, negotiative alignment systematically identifies persistent conflicts, assesses whether they can be partly resolved through design modifications, and updates stakeholder-specific weighting factors to reflect each group’s preferences. By embedding formal bargaining mechanisms (Nash, 1950; Kalai and Smorodinsky, 1975) in the design pipeline, negotiative alignment aims to enhance fairness and inclusivity. Far from viewing divergence as an anomaly, it treats disagreement as a resource for building more equitable outcomes.
1.1. Contribution and Research Questions
@@
  • (RQ2) How do different evaluation methods capture disagreement, particularly among minority groups with intersectional identities? While ranking and rating methods are well-established in social choice theory for aggregating preferences (Leroy et al., 2024; Xu et al., 2024; Rofin et al., 2023; Suzuki and Horita, 2024; Fritz et al., 2023), we seek to understand how these methods reflect the agreement and disagreement dynamics within minority populations possessing intersectional identities. This focus allows us to examine whether traditional evaluation techniques adequately capture the nuanced preferences and potential conflicts that arise from overlapping social categories such as race, gender, and socioeconomic status.
@@ 2. Related Literature
Urban assessment frameworks commonly aggregate feedback into a single metric for decision-making (Mehta, 2014; Varna and Tiesdell, 2010; Naik et al., 2014; Qian et al., 2023; Yurrita et al., 2021). Although efficient for policy or administrative planning, these techniques do not always capture the nuanced needs of marginalized communities (Fainstein, 2010; Sandercock, 1998; Gowaikar et al., 2024). Inaccessible sidewalks (Imrie, 2012) and queer space erasure (Doan, 2011) exemplify how aggregate metrics may obscure issues relevant to minority groups. Our findings align with these critiques, suggesting that demographic-specific experiences can yield systematic disagreements worthy of deeper consideration.
@@
Urban planners have historically relied on frameworks such as the Gehl Institute’s public life tools (Gehl and Svarre, 2013) and the “Place Diagram” by Project for Public Spaces (Santos Nouri and Costa, 2017), as well as indices like the Public Space Index (PSI) (Mehta, 2014) and the Pedestrian Environment Quality Index (PSEQI) (Zamanifard et al., 2019) to evaluate aspects of comfort, access, and sociability. While these approaches incorporate observational data and community engagement, they typically yield composite metrics that may mask diverse and context-specific experiences. Recent scholarship in critical urban studies highlights the limitations of consensus-based measures, particularly their potential to underrepresent the needs of marginalized groups. Our work extends this critique by proposing a framework that employs participatory methods to formalize disagreement through a negotiative lens, thereby complementing established approaches and advancing discussions on equity in urban design.
2.1. AI Alignment and Pluralistic Models
@@
Traditional AI alignment generally involves mapping a model’s objective function to a representative user preference or a set of aggregated preferences (Gabriel, 2020; Rahwan et al., 2019). This approach works adequately for many applications, but it often overlooks persistent disagreements and minority voices. Pluralistic alignment methods (Zhi-Xuan et al., 2024; Sorensen et al., 2024; Nayak et al., 2024) have begun to address this limitation by modeling a distribution of perspectives rather than a single consensus. However, these techniques usually stop short of incorporating active negotiation or formal bargaining procedures to handle direct conflicts between stakeholder groups.
2.1.1. Multi-Agent Reinforcement Learning and Game Theory
@@
Multi-Agent Reinforcement Learning (MARL) extends single-agent paradigms by assigning separate objectives to each agent, sometimes using equilibrium concepts or Pareto-efficiency to mediate conflicts (Zhang et al., 2021; Yang et al., 2020). Game-theoretic approaches, including Nash bargaining (Nash, 1950; Nakamura, 2024), Kalai-Smorodinsky fairness (Kalai and Smorodinsky, 1975; Chen and Hooker, 2023), and cooperative resource allocation (Rosen, 1965; Doostmohammadian et al., 2025), also provide structured ways to reconcile divergent interests. While these models facilitate coordination among agents with different goals, they typically assume a static set of fairness constraints and do not dynamically adjust stakeholder influence as new information or disagreements arise.
2.1.2. Deliberative AI and Fairness-Conscious Optimization
@@
Deliberative AI focuses on argumentation and structured dialogue to find acceptable outcomes (Walton et al., 2009; Rahwan and Simari, 2009; Nakamura, 2024; Fish et al., 2023). Although such approaches emphasize reasoning about conflicting arguments, they often lack iterative mechanisms to update weights across multiple negotiation rounds. Similarly, fairness-conscious optimization seeks to maintain acceptable performance across demographic groups (Moulin, 2004; Dwork et al., 2012; Nakamura, 2024), but typically imposes static fairness constraints. Disagreement persists if preferences evolve or if new stakeholder groups become relevant.
2.2. Negotiative Alignment
@@
To handle disagreements that neither fade over time nor converge under static fairness criteria, negotiative alignment incorporates dynamic bargaining mechanisms. This extension is particularly relevant in urban planning, where stakeholder needs can shift due to evolving infrastructure, demographics, or cultural contexts. Past research in multi-agent urban planning simulates stakeholders with distinct objectives (Qian et al., 2023; Yurrita et al., 2021; Zhou et al., 2024; Ni et al., 2024), but it rarely integrates formal negotiation procedures to address deep-seated conflicts. Dissent-aware learning methods likewise highlight the importance of preserving minority viewpoints (Gordon et al., 2022; Chen et al., 2021), yet they do not specify how to arbitrate among directly opposing interests once disagreements are identified.
@@
3. Methodology
3.1. Study Setting: Montreal, Summer 2023–Autumn 2024
We conducted this study in Montreal, a multilingual city characterized by diverse architectural styles and intersecting social communities. This diversity enabled recruitment from multiple marginalized populations, thereby illustrating the variety of urban experiences and how conflicting preferences necessitate a framework for negotiative alignment (Margier, 2013; Sylvestre, 2010; Mushkani et al., 2025b).
@@ 3.1. Study Setting: Montreal, Summer 2023–Autumn 2024
Figure 1. Representation of diversity across age groups. The distribution of self-reported demographic characteristics varies by age, highlighting intersectional perspectives and reminding us that no single identity alone defines an individual’s experience (Mushkani et al., 2025b).
[…]
Representation of diversity across age groups
3.2. Participant Recruitment and Profiles
We contacted over 100 local organizations and established partnerships with more than 30 institutions serving wheelchair users, seniors, LGBTQIA2+ individuals, and recent immigrants (IRCGM, 2018; Mushkani et al., 2025b). Although including a majority-resident control group might have offered broader contrasts, our focus was on systematically marginalized participants, whose perspectives are often obscured by aggregate metrics. As a result of these partnerships, 35 individuals participated in the study, with several contributing to multiple phases of data collection:
@@ 3.2. Participant Recruitment and Profiles
3.3. Multi-Phase Data Collection
3.3.1. Phase I: Semi-Structured Interviews
We conducted 28 semi-structured interviews focusing on participants’ conceptualizations of quality in public spaces. Each interview lasted between 40 and 90 minutes and was audio recorded. Questions included how participants use public spaces, how they define and talk about them, which ones are their favorites, and what kinds of experiences they recall when discussing an ideal public space. Additionally, we explored what qualities participants believe an ideal public space should possess. These initial questions were followed by more contextual and in-depth inquiries (Creswell and Creswell, 2022).
3.3.2. Phase II: Focus Groups and Rating
Three focus-group sessions, each consisting of 3–5 participants, were organized based on participants’ availability and schedules. During these sessions, individuals first independently rated 20 selected street images on a 4-point scale (1,=,poor; 4,=,excellent) according to four core criteria: inclusivity, accessibility, aesthetics, and practicality. Our selection of 20 street views adhered to a structured socio-spatial matrix (see Appendix Figures 11–13), ensuring representational diversity in land-use, historical contexts, and green-space distribution. Following the individual ratings, participants engaged in group discussions to negotiate their ratings, exploring whether disagreements were due to objective observations (e.g., unrecognized curb ramps) or deeper value differences (e.g., differing notions of safety) (Margier, 2013).
@@ 3.3.2. Phase II: Focus Groups and Rating
3.3.3. Phase III: Ranking Experiment
Seventeen participants were tasked with individually ranking seven images from best to worst based on each of the 12 previously identified criteria. This process resulted in twelve sets of rankings per participant, each corresponding to a different criterion. For each set, Kendall’s Tau was computed to measure the level of agreement among participants regarding the positioning of the images. The ranking exercise compelled participants to make explicit trade-offs, thereby revealing underlying priorities and values influencing their evaluations of urban spaces (Sylvestre, 2010). Due to the nature of this task, we reduced the number of images from 20 to 7, allowing us to increase the number of criteria and compare the performance of various concepts more effectively.
Our partnership with community organizations remains ongoing, emphasizing a co-production philosophy through workshops and knowledge exchange. Participants were compensated for their time and contributions, while we also fostered a reciprocal relationship in which participants learned about AI and expert knowledge on cities from the researchers, while we gained insights into public space usage and perceptions from the citizens (IRCGM, 2018; Mushkani et al., 2025a). Furthermore, all study procedures were reviewed and approved by the Research Ethics Board, ensuring that ethical standards were maintained throughout the research process.
3.4. Analysis Methods
@@
We thematically coded interview and focus-group transcripts, emphasizing statements that elucidated the underlying why of disagreements (e.g., cultural symbolism, mobility constraints) (Creswell and Creswell, 2022). From the interviews, we identified twelve criteria that were used for the ranking task: accessibility, invitingness, comfort, regenerativeness, beauty, practicality, maintenance, inclusivity, dynamism, representativeness, oppression, and security.
@@
  • Pearson’s Correlation (Puth et al., 2015): We calculated agreement levels between individual and group ratings, as well as between different demographic groups. This analysis assessed the extent to which group discussions influenced rating shifts, identifying cases where collective deliberation significantly altered individual assessments.
  • Kendall’s Tau (Abdi, 2006): We measured ordinal ranking similarities across demographic groups for each criterion to identify patterns of agreement and areas of persistent disagreement. This helped pinpoint criteria where differences in weighting were irreconcilable, indicating fundamental divergences in preferences.
@@
4. Results
4.1. Overview of Street-Level Findings

### v2

@@ Abstract
Based on these empirical insights, we propose negotiative alignment, an AI framework that treats disagreement as an essential input to be preserved, analyzed, and addressed. Negotiative alignment builds on pluralistic models by dynamically updating stakeholder preferences through multi-agent negotiation mechanisms, ensuring no single perspective is marginalized. We outline how this framework can be integrated into urban analytics—and other decision-making contexts—to retain minority viewpoints, adapt to changing stakeholder concerns, and enhance fairness and accountability. The study demonstrates that preserving and engaging with disagreement, rather than striving for an artificial consensus, can produce more equitable and responsive outcomes in urban design.
@@
Keywords: Negotiative alignment, urban design, AI alignment, multi-stakeholder preferences, fairness, accountability, participatory, inclusivity
Introduction
@@
Urban environments appear to serve everyone, yet people experience cities through different social, cultural, and political lenses. A wheelchair user in Montreal may perceive certain sidewalks as insurmountable, while an LGBTQIA2+ resident might interpret street murals or local businesses as signaling inclusion or exclusion. However, contemporary urban analysis tends to emphasize aggregated metrics: a single “overall rating” for aesthetic quality, accessibility, or safety (Mehta 2014; Varna and Tiesdell 2010). Although helpful for city-wide planning, these averages may dilute the specific, and often urgent, needs of marginalized groups (Fainstein 2010; Costanza-Chock 2020).
@@
Drawing on a multi-phase study of public space in Montreal, we show that apparent disagreements in urban evaluations are not merely noisy outliers. Rather, they reflect differences in lived experience, structural inequality, and cultural values. Participants in our study included 35 individuals from multiple demographic categories, such as wheelchair users, seniors, LGBTQIA2+ individuals, and immigrants. By combining semi-structured interviews, rating tasks, and forced-ranking experiments, we found that disagreements were pervasive and persistent, particularly concerning dimensions like accessibility, inclusivity, and practicality. As discussed in Section Results, such divergent viewpoints could not be resolved merely by factual clarifications; instead, they pointed to deeper identity-based needs.
@@
Conventional AI or survey-based methods prioritize consensus by aggregating heterogeneous preferences into a single output score. This approach often obscures critical minority viewpoints (Gabriel 2020; Rawls 1999), such as the need for comprehensive curb cuts for wheelchair users or the importance of affirming cultural symbols for LGBTQIA2+ communities (Anttiroiko and de Jong 2020). While recent advances in AI alignment incorporate diverse perspectives through multi-agent modeling, game-theoretic approaches, and fairness-aware optimization (Zhi-Xuan et al. 2024; Sorensen et al. 2024; Jain et al. 2024; Dwork et al. 2012; Ferrara 2023), they rarely include explicit mechanisms for negotiating conflicting preferences in real time. Addressing this gap requires alignment strategies that iteratively adjust stakeholder weights and trade-offs rather than statically aggregating preferences.
@@
To bridge this gap, we introduce negotiative alignment, a dynamic, multi-agent approach that integrates disagreements into iterative AI decision-making processes (Section Negotiative Alignment Framework). Unlike singular consensus methods, negotiative alignment systematically identifies persistent conflicts, assesses whether they can be partly resolved through design modifications, and updates stakeholder-specific weighting factors to reflect each group’s preferences. By embedding formal bargaining mechanisms (Nash 1950; Kalai and Smorodinsky 1975) in the design pipeline, negotiative alignment aims to enhance fairness and inclusivity. Far from viewing divergence as an anomaly, it treats disagreement as a resource for building more equitable outcomes.
Contribution and Research Questions
@@
  • (RQ2) How do different evaluation methods capture disagreement, particularly among minority groups with intersectional identities? While ranking and rating methods are well-established in social choice theory for aggregating preferences (Leroy et al. 2024; Xu et al. 2024; Rofin et al. 2023; Suzuki and Horita 2024; Fritz, Moretti, and Staudacher 2023), we seek to understand how these methods reflect the agreement and disagreement dynamics within minority populations possessing intersectional identities. This focus allows us to examine whether traditional evaluation techniques adequately capture the nuanced preferences and potential conflicts that arise from overlapping social categories such as race, gender, and socioeconomic status.
@@
Urban assessment frameworks commonly aggregate feedback into a single metric for decision-making (Mehta 2014; Varna and Tiesdell 2010; Naik et al. 2014; Qian et al. 2023; Yurrita et al. 2021). Although efficient for policy or administrative planning, these techniques do not always capture the nuanced needs of marginalized communities (Fainstein 2010; Sandercock 1998; Gowaikar et al. 2024). Inaccessible sidewalks (Imrie 2012) and queer space erasure (Doan 2011) exemplify how aggregate metrics may obscure issues relevant to minority groups. Our findings align with these critiques, suggesting that demographic-specific experiences can yield systematic disagreements worthy of deeper consideration.
@@
Urban planners have historically relied on frameworks such as the Gehl Institute’s public life tools (Gehl and Svarre 2013) and the “Place Diagram” by Project for Public Spaces (Santos Nouri and Costa 2017), as well as indices like the Public Space Index (PSI) (Mehta 2014) and the Pedestrian Environment Quality Index (PSEQI) (Zamanifard et al. 2019) to evaluate aspects of comfort, access, and sociability. While these approaches incorporate observational data and community engagement, they typically yield composite metrics that may mask diverse and context-specific experiences. Recent scholarship in critical urban studies highlights the limitations of consensus-based measures, particularly their potential to underrepresent the needs of marginalized groups. Our work extends this critique by proposing a framework that employs participatory methods to formalize disagreement through a negotiative lens, thereby complementing established approaches and advancing discussions on equity in urban design.
AI Alignment and Pluralistic Models
@@
Traditional AI alignment generally involves mapping a model’s objective function to a representative user preference or a set of aggregated preferences (Gabriel 2020; Rahwan et al. 2019). This approach works adequately for many applications, but it often overlooks persistent disagreements and minority voices. Pluralistic alignment methods (Zhi-Xuan et al. 2024; Sorensen et al. 2024; Nayak et al. 2024) have begun to address this limitation by modeling a distribution of perspectives rather than a single consensus. However, these techniques usually stop short of incorporating active negotiation or formal bargaining procedures to handle direct conflicts between stakeholder groups.
Multi-Agent Reinforcement Learning and Game Theory
@@
Multi-Agent Reinforcement Learning (MARL) extends single-agent paradigms by assigning separate objectives to each agent, sometimes using equilibrium concepts or Pareto-efficiency to mediate conflicts (Zhang, Yang, and Başar 2021; Yang et al. 2020). Game-theoretic approaches, including Nash bargaining (Nash 1950; Nakamura 2024), Kalai-Smorodinsky fairness (Kalai and Smorodinsky 1975; Chen and Hooker 2023), and cooperative resource allocation (Rosen 1965; Doostmohammadian et al. 2025), also provide structured ways to reconcile divergent interests. While these models facilitate coordination among agents with different goals, they typically assume a static set of fairness constraints and do not dynamically adjust stakeholder influence as new information or disagreements arise.
Deliberative AI and Fairness-Conscious Optimization
@@
Deliberative AI focuses on argumentation and structured dialogue to find acceptable outcomes (Walton, Reed, and Macagno 2009; Rahwan and Simari 2009; Nakamura 2024; Fish et al. 2023). Although such approaches emphasize reasoning about conflicting arguments, they often lack iterative mechanisms to update weights across multiple negotiation rounds. Similarly, fairness-conscious optimization seeks to maintain acceptable performance across demographic groups (Moulin 2004; Dwork et al. 2012; Nakamura 2024), but typically imposes static fairness constraints. Disagreement persists if preferences evolve or if new stakeholder groups become relevant.
Negotiative Alignment
@@
To handle disagreements that neither fade over time nor converge under static fairness criteria, negotiative alignment incorporates dynamic bargaining mechanisms. This extension is particularly relevant in urban planning, where stakeholder needs can shift due to evolving infrastructure, demographics, or cultural contexts. Past research in multi-agent urban planning simulates stakeholders with distinct objectives (Qian et al. 2023; Yurrita et al. 2021; Zhou et al. 2024; Ni, Wang, and Liu 2024), but it rarely integrates formal negotiation procedures to address deep-seated conflicts. Dissent-aware learning methods likewise highlight the importance of preserving minority viewpoints (Gordon et al. 2022; Chen, Weld, and Zhang 2021), yet they do not specify how to arbitrate among directly opposing interests once disagreements are identified.
@@
Methodology
Study Setting: Montreal, Summer 2023–Autumn 2024
We conducted this study in Montreal, a multilingual city characterized by diverse architectural styles and intersecting social communities. This diversity enabled recruitment from multiple marginalized populations, thereby illustrating the variety of urban experiences and how conflicting preferences necessitate a framework for negotiative alignment (Margier 2013; Sylvestre 2010; Mushkani et al. 2025b).
@@
Figure 1: Representation of diversity across age groups. The distribution of self-reported demographic characteristics varies by age, highlighting intersectional perspectives and reminding us that no single identity alone defines an individual’s experience (Mushkani et al. 2025b).
Participant Recruitment and Profiles
We contacted over 100 local organizations and established partnerships with more than 30 institutions serving wheelchair users, seniors, LGBTQIA2+ individuals, and recent immigrants (IRCGM 2018; Mushkani et al. 2025b). Although including a majority-resident control group might have offered broader contrasts, our focus was on systematically marginalized participants, whose perspectives are often obscured by aggregate metrics. As a result of these partnerships, 35 individuals participated in the study, with several contributing to multiple phases of data collection:
@@
Multi-Phase Data Collection
Phase I: Semi-Structured Interviews
We conducted 28 semi-structured interviews focusing on participants’ conceptualizations of quality in public spaces. Each interview lasted between 40 and 90 minutes and was audio recorded. Questions included how participants use public spaces, how they define and talk about them, which ones are their favorites, and what kinds of experiences they recall when discussing an ideal public space. Additionally, we explored what qualities participants believe an ideal public space should possess. These initial questions were followed by more contextual and in-depth inquiries (Creswell and Creswell 2022).
Phase II: Focus Groups and Rating
Three focus-group sessions, each consisting of 3–5 participants, were organized based on participants’ availability and schedules. During these sessions, individuals first independently rated 20 selected street images on a 4-point scale (1,=,poor; 4,=,excellent) according to four core criteria: inclusivity, accessibility, aesthetics, and practicality. Our selection of 20 street views adhered to a structured socio-spatial matrix (see Appendix Figures 11–13), ensuring representational diversity in land-use, historical contexts, and green-space distribution. Following the individual ratings, participants engaged in group discussions to negotiate their ratings, exploring whether disagreements were due to objective observations (e.g., unrecognized curb ramps) or deeper value differences (e.g., differing notions of safety) (Margier 2013).
@@
Phase III: Ranking Experiment
Seventeen participants were tasked with individually ranking seven images from best to worst based on each of the 12 previously identified criteria. This process resulted in twelve sets of rankings per participant, each corresponding to a different criterion. For each set, Kendall’s Tau was computed to measure the level of agreement among participants regarding the positioning of the images. The ranking exercise compelled participants to make explicit trade-offs, thereby revealing underlying priorities and values influencing their evaluations of urban spaces (Sylvestre 2010). Due to the nature of this task, we reduced the number of images from 20 to 7, allowing us to increase the number of criteria and compare the performance of various concepts more effectively.
Our partnership with community organizations remains ongoing, emphasizing a co-production philosophy through workshops and knowledge exchange. Participants were compensated for their time and contributions, while we also fostered a reciprocal relationship in which participants learned about AI and expert knowledge on cities from the researchers, while we gained insights into public space usage and perceptions from the citizens (IRCGM 2018; Mushkani et al. 2025a). Furthermore, all study procedures were reviewed and approved by the Research Ethics Board, ensuring that ethical standards were maintained throughout the research process.
Analysis Methods
@@
We thematically coded interview and focus-group transcripts, emphasizing statements that elucidated the underlying why of disagreements (e.g., cultural symbolism, mobility constraints) (Creswell and Creswell 2022). From the interviews, we identified twelve criteria that were used for the ranking task: accessibility, invitingness, comfort, regenerativeness, beauty, practicality, maintenance, inclusivity, dynamism, representativeness, oppression, and security.
@@
  • Pearson’s Correlation (Puth, Neuhäuser, and Ruxton 2015): We calculated agreement levels between individual and group ratings, as well as between different demographic groups. This analysis assessed the extent to which group discussions influenced rating shifts, identifying cases where collective deliberation significantly altered individual assessments.
  • Kendall’s Tau (Abdi 2006): We measured ordinal ranking similarities across demographic groups for each criterion to identify patterns of agreement and areas of persistent disagreement. This helped pinpoint criteria where differences in weighting were irreconcilable, indicating fundamental divergences in preferences.
@@
Results
Overview of Street-Level Findings